# Low-temperature scaling laws in unconventional flat-band superconductors

### v1

@@ I Introduction
When ignoring interband pairing (which gives rise to offsets in the scaling laws of the density of states determined by the strength of a pseudo-magnetic field [16]), the low-temperature scaling laws of observables are completely fixed by the topology of the zeros, i.e. nodes, of the gap function in momentum space [17, 10]. One key quantity that indicates the possibility of superconductivity is the superfluid weight (or superfluid stiffness) [18]. In the absence of Galilean invariance, it has been shown that nontrivial single-particle quantum geometry characterizes the underlying mechanism responsible for the existence of a nonzero superfluid weight in a flat band with divergent effective mass [19, 20]. While it is well known that the “minimal quantum metric” is solely responsible for a nonzero superfluid weight in conventional superconductors with isolated flat bands [21], recent studies indicate the appearance of nonlocal quantum-geometric terms in the superfluid weight for the majority of unconventional pairing due to the momentum-dependent nature of the gap function [20, 22, 23]. Therefore, it is important to clarify whether the scaling laws obtained in Ref. [24] for the superfluid weight of flat-band superconductors need to be adjusted when taking into account the arising functional, i.e. nonlocal quantum geometrical, superfluid weight.
For completeness, we additionally calculate the low-temperature scaling laws of the order parameter, the tunneling conductance, the specific heat, the Sommerfeld coefficient, and the spin-lattice relaxation rate for flat-band superconductors with nontrivial nodal structure, providing a full guide for experimental measurements. Our main results are summarized in Table 1. See also Ref. [25], in which dynamical mean-field theory calculations for an attractive Hubbard model on the Lieb lattice indicate that the superfluid weight follows a Gorter-Casimir-like behavior [26, 27].
Table 1: Collection of low-temperature scaling laws in flat-band superconductors with different nodal structures covered by the dispersion (18) for the order parameter $\Delta_{T}$, the geometrical and functional superfluid weights $D_{\mathrm{s}}^{\mathrm{geom},\mathrm{func}}$, the tunneling conductance $G_{sn}$, the specific heat $C$, the Sommerfeld coefficient $\gamma$, and the NMR spin-lattice relaxation rate $1/(T_{1}T)$. Here, $m>0$ represents the total order of vanishing of the gap function and $L$ indicates the number of straight nodal lines through the origin, i.e., the case of $L=1$ describes a single line node without crossings, $L=2$ a crossing of two line nodes, and $L>2$ crossings of three or more line nodes.
@@
point node $T^{\frac{2}{m}+1}$ $T^{\frac{2}{m}+1}$ $T^{\frac{4}{m}+2}$ $T^{\frac{2}{m}-1}$ $T^{\frac{2}{m}}$ $T^{\frac{2}{m}-1}$ $T^{\frac{4}{m}-2}$
line node ($L=1$) $T^{\frac{1}{m}+1}$ $T^{\frac{1}{m}+1}$ $T^{\frac{2}{m}+2}$ $T^{\frac{1}{m}-1}$ $T^{\frac{1}{m}}$ $T^{\frac{1}{m}-1}$ $T^{\frac{2}{m}-2}$
line node ($L=2$) $T^{\frac{2}{m}+1}\ln(1/T)$ $T^{\frac{2}{m}+1}\ln(1/T)$ $T^{\frac{4}{m}+2}\ln^{2}(1/T)$ $T^{\frac{2}{m}-1}\ln(1/T)$ $T^{\frac{2}{m}}\ln(1/T)$ $T^{\frac{2}{m}-1}\ln(1/T)$ $T^{\frac{4}{m}-2}\ln^{2}(1/T)$
line node ($L>2$) $T^{\frac{2}{m}+1}$ $T^{\frac{2}{m}+1}$ $T^{\frac{4}{m}+2}$ $T^{\frac{2}{m}-1}$ $T^{\frac{2}{m}}$ $T^{\frac{2}{m}-1}$ $T^{\frac{4}{m}-2}$
@@ II Nodal structure of the gap function
The grand potential of mean-field BCS theory for nonzero temperature is given by [31, 21, 23]
@@
$T$ represents the temperature, $\varepsilon_{\mathbf{k}}=\mathrm{diag}(\varepsilon_{\mathbf{k}1},\ldots,\varepsilon_{\mathbf{k}N_{\mathrm{B}}})$ contains the $N_{\mathrm{B}}$ eigenvalues of the time-reversal symmetric (TRS) single-particle Hamiltonian $H(\mathbf{k})$, $\mu$ is the chemical potential, $U(\mathbf{k},\mathbf{k}^{\prime})$ is the effective pairing potential, and $\Delta_{\alpha\beta}(\mathbf{q};\mathbf{k})$ is the gap function with band indices $\alpha,\beta=1,\ldots,N_{\mathrm{B}}$ which contains the order parameter. Note that the formalism of this work allows the consideration of TRS-breaking pairing mechanisms.
The classification of superconducting states is done via the irreducible representations of the symmetry group of the system [32]. In particular, we need to distinguish between spin-singlets with total spin $S=0$ and spin-triplets with total spin $S=1$. The pair spin wave function of a spin-singlet is antisymmetric with respect to an exchange of the spin indices, i.e. the gap function can be expressed as
@@
where the function $f_{\alpha\beta}$ is determined by the symmetry group of the system. If $\Gamma$ denotes an irreducible representation of the group with dimension $d_{\Gamma}$, it is provided by
@@
where $\psi_{i}^{\Gamma}$ are the simplest basis functions that are even in $\mathbf{k}$ and respect the symmetry of the system and $\Delta_{\alpha\beta}^{i}$ are coefficients representing the order parameters of the superconductor. Similarly, a spin-triplet state with total spin $S=1$ has odd parity, i.e. the gap function can be expressed as
@@
$\Delta_{\alpha\beta}^{i}=\frac{U_{0}}{V}\sum_{\mathbf{k},n}\bar{\psi}_{i}^{\Gamma}(\mathbf{k})v_{n\alpha,+}(\mathbf{k})v^{\ast}_{n\beta,-}(\mathbf{k})(1-2n_{\mathrm{F}}(E_{\mathbf{k}n}))$ (8)
at $\mathbf{q}=0$. Here, the $2N_{\mathrm{B}}$ vectors $(v_{n,+},v_{n,-})$ represent the eigenvectors of the BdG Hamiltonian in the Nambu spinor basis. Moreover, we assume for simplicity that the superconducting state is fully characterized by one order parameter $\Delta_{T}$ which represents the $\mathbf{k}$-independent but temperature-dependent proportionality factor of the gap function above, i.e. $\Delta_{\alpha\beta}^{i}\equiv\Delta_{T}\delta_{\alpha\beta}$. Analogously to Refs. [10, 32, 33], the self-consistent equation can be written as
$\Delta_{T}=\frac{U_{0}}{2V}\sum_{\mathbf{k},n}\frac{\bar{\psi}_{i}^{\Gamma}(\mathbf{k})f(\mathbf{k})}{E_{\mathbf{k}n}}(1-2n_{\mathrm{F}}(E_{\mathbf{k}n}))\,.$ (9)
The quasiparticle dispersion in a flat band can be approximated by $E_{\mathbf{k}}\approx|f(\mathbf{k})|$, i.e., the nodal structure of a flat-band superconductor is dictated by the gap function. Let us shift and rotate the coordinate system such that the analytic function $f$ has a node at the origin and is not identically zero along the $k_{1}$-direction. The Weierstrass preparation theorem states that there exists a Weierstrass polynomial,
@@
As is clear from the factorization, the zero set of each linear factor with $\mathrm{Im}(\lambda_{i})=0$ produces a line node $k_{1}-\mathrm{Re}\,\lambda_{i}(k_{2})=0$ with shallowness $m_{i}$, while it becomes a point node of shallowness $m_{i}$ as soon as $\mathrm{Im}(\lambda_{i})\neq 0$. To distinguish both types, we also write the factorization as
@@
exhibits a “cusp” at the origin. Nevertheless, the Newton-Puiseux theorem guarantees here that the functions $u_{i},v_{i},w_{i}$ admit convergent Puiseux expansions [35].
@@
Because the calculations become very complicated if one considers the general expression for the dispersion given in Eq. (15), this work restricts ourselves to the representative case with $v_{1,\ldots,J}\equiv 0$, $w_{1,\ldots,J}(k_{2})=k_{2}$, and
@@
where $m=qL+\sum_{j=1}^{J}p_{j}$ is the total radial order of vanishing, $L$ is associated to the number of straight nodal lines through the origin, and $q$ is the shallowness per line, i.e. $E_{\mathbf{k}}\sim\Delta_{T}|t|^{q}$ near any of the nodal lines if local coordinates with $t$ transverse to the line are taken. Moreover, to simplify the calculations, we assume $J=0$ if $L>0$, so that the pointlike enhancement at the origin is absent and $m=qL$ for the line nodes. This is enough for our physical application later.
@@ III.2 Line nodes
$D(E)=\frac{1}{4\pi^{2}}\int_{0}^{2\pi}\mathrm{d}\theta\int_{0}^{\Lambda}\mathrm{d}k\,k\,\delta(E-k^{m}|\cos(L\theta)|^{q})\,.$ (24)
@@
The incomplete beta function reads [37, §8.39]
@@
which for $x=1$ corresponds to the usual beta function. Therefore, we find the DOS to be given by
@@
For $q/m=1/2$, we make use of the fact that $B_{x}(1/2,0)=2\mathrm{artanh}(\sqrt{x})$ [37, §9.12]. Since we have $\mathrm{artanh}(\sqrt{1-x^{2}})=\ln(2/x)+\mathcal{O}(x^{2})$ for $x\ll 1$, we find at leading order
@@
We proceed similarly for $q/m>1/2$. By using the power series of the hypergeometric function [37, §9.14], we find
@@
with $x=(E/(\Delta_{T}\Lambda^{m}))^{2/q}$ in our case. Thus, we obtain the renormalized DOS as
@@ IV.1 Order parameter
$d_{\Gamma}\Delta_{T}\approx\frac{U_{0}}{2V}\sum_{\mathbf{k}}\frac{|f(\mathbf{k})|^{2}}{\Delta_{T}E_{\mathbf{k}}}(1-2n_{\mathrm{F}}(E_{\mathbf{k}}))\,$, (34)
where we inserted $f(\mathbf{k})=\Delta_{T}\sum_{i=1}^{d_{\Gamma}}\psi_{i}^{\Gamma}(\mathbf{k})$, cf. Eq. (4). We further replace the momentum integral by an energy integral via Eq. (19). The gap equation becomes
$\Delta_{T}=\left[\frac{U_{0}}{2d_{\Gamma}}\int_{0}^{\infty}\mathrm{d}E\,D(E)E(1-2n_{\mathrm{F}}(E))\right]^{1/2}\,.$ (35)
@@
$\Delta_{T}=\left[\frac{U_{0}\tilde{D}_{0}}{2d_{\Gamma}}\int_{0}^{\infty}\mathrm{d}E\,E^{\alpha}\ln^{\beta}(1/E)(1-2n_{\mathrm{F}}(E))\right]^{1/(2+\alpha)}\,.$ (37)
@@
$\Delta_{0}-\Delta_{T}=\frac{U_{0}\tilde{D}_{0}}{d_{\Gamma}(\alpha+2)\Delta_{0}^{\alpha+1}}\int_{0}^{\infty}\mathrm{d}E\,\frac{E^{\alpha}\ln^{\beta}(1/E)}{e^{E/T}+1}\,.$ (38)
@@
$C_{\alpha}=\Gamma(\alpha+1)\eta(\alpha+1)\frac{U_{0}\tilde{D}_{0}}{d_{\Gamma}(\alpha+2)\Delta_{0}^{\alpha+1}}\,.$ (40)

### v2

@@ I Introduction
When ignoring interband pairing (which gives rise to offsets in the scaling laws of the density of states determined by the strength of a pseudo-magnetic field [30]), the low-temperature scaling laws of observables are completely fixed by the topology of the zeros, i.e., nodes, of the gap function in momentum space [53, 46]. One key quantity that indicates the possibility of superconductivity is the superfluid weight (or superfluid stiffness) [32]. In the absence of Galilean invariance, it has been shown that nontrivial single-particle quantum geometry characterizes the underlying mechanism responsible for the existence of a nonzero superfluid weight in a flat band with divergent effective mass [24, 60]. While the minimal quantum metric is solely responsible for a nonzero superfluid weight in conventional superconductors with isolated flat bands [22], recent studies indicate the appearance of nonlocal quantum-geometric terms in the superfluid weight for the majority of unconventional pairing due to the momentum-dependent nature of the gap function [60, 29, 9]. Therefore, it is important to clarify whether the scaling laws obtained in Ref. [20] for the superfluid weight of flat-band superconductors need to be adjusted when taking into account the arising functional, i.e., nonlocal quantum geometrical, superfluid weight.
For completeness, we additionally calculate the low-temperature scaling laws of the order parameter, the tunneling conductance, the specific heat, the Sommerfeld coefficient, and the spin-lattice relaxation rate for flat-band superconductors with nontrivial nodal structure, providing a full guide for experimental measurements. Our main results are summarized in Table 1. See also Ref. [44], where the authors discuss the low-temperature behavior of the superfluid weight in twisted graphene in the presence of band-off-diagonal pairing, and Ref. [40], in which dynamical mean-field theory calculations for an attractive Hubbard model on the Lieb lattice indicate that the superfluid weight follows a Gorter-Casimir-like behavior [51, 41].
Table 1: Collection of low-temperature scaling laws in flat-band superconductors with different nodal structures covered by the dispersion (18) for the order parameter $\Delta_{T}$, the geometrical and functional superfluid weights $D_{\mathrm{s}}^{\mathrm{geom},\mathrm{func}}$, the tunneling conductance $G_{sn}$, the specific heat $C$, the Sommerfeld coefficient $\gamma$, and the NMR spin-lattice relaxation rate $1/(T_{1}T)$. Here, $m>0$ represents the total order of vanishing of the gap function and $L$ indicates the number of straight nodal lines through the origin, i.e., the case of $L=1$ describes a single line node without crossings, $L=2$ a crossing of two line nodes, and $L>2$ crossings of three or more line nodes. Moreover, $b\geq 0$ represents a superconducting state dependent exponent which is determined by Eq. (69). We have $b=1$ for all one-dimensional superconducting states as well as for all two-dimensional states considered in Sec. V.
@@
point node $T^{\frac{2}{m}+1}$ $T^{\frac{2}{m}+1}$ $T^{{\frac{2}{m}+b}}$ $T^{\frac{2}{m}-1}$ $T^{\frac{2}{m}}$ $T^{\frac{2}{m}-1}$ $T^{\frac{4}{m}-2}$
line node ($L=1$) $T^{\frac{1}{m}+1}$ $T^{\frac{1}{m}+1}$ $T^{{\frac{1}{m}+b}}$ $T^{\frac{1}{m}-1}$ $T^{\frac{1}{m}}$ $T^{\frac{1}{m}-1}$ $T^{\frac{2}{m}-2}$
line node ($L=2$) $T^{\frac{2}{m}+1}\ln(1/T)$ $T^{\frac{2}{m}+1}\ln(1/T)$ $T^{{\frac{2}{m}+b}}\ln(1/T)$ $T^{\frac{2}{m}-1}\ln(1/T)$ $T^{\frac{2}{m}}\ln(1/T)$ $T^{\frac{2}{m}-1}\ln(1/T)$ $T^{\frac{4}{m}-2}\ln^{2}(1/T)$
line node ($L>2$) $T^{\frac{2}{m}+1}$ $T^{\frac{2}{m}+1}$ $T^{{\frac{2}{m}+b}}$ $T^{\frac{2}{m}-1}$ $T^{\frac{2}{m}}$ $T^{\frac{2}{m}-1}$ $T^{\frac{4}{m}-2}$
@@ II Nodal structure of the gap function
The grand potential of mean-field BCS theory for nonzero temperature is given by [22]
@@
$T$ represents the temperature, $\varepsilon_{\mathbf{k}}=\mathrm{diag}(\varepsilon_{\mathbf{k}1},\ldots,\varepsilon_{\mathbf{k}N_{\mathrm{B}}})$ contains the $N_{\mathrm{B}}$ eigenvalues of the time-reversal symmetric (TRS) single-particle Hamiltonian $H(\mathbf{k})$, $\mu$ is the chemical potential, $U(\mathbf{k},\mathbf{k}^{\prime})$ is the complex-valued effective pairing potential, and $\Delta_{\alpha\beta}(\mathbf{q};\mathbf{k})$ is the gap function with band indices $\alpha,\beta=1,\ldots,N_{\mathrm{B}}$ which contains the order parameter. Note that the formalism of this work allows the consideration of TRS-breaking pairing mechanisms, but not TRS-breaking single-particle Hamiltonians. If the single-particle Hamiltonian $H(\mathbf{k})$ exhibits a flat band, its time-reversed partner has the same spectrum (up to $\mathbf{k}\to-\mathbf{k}$), and hence, exhibits the same flat band. Consequently, the DOS and the corresponding scaling laws of the order parameter, the tunneling conductance, the specific heat, the Sommerfeld coefficient, and the NMR relaxation rate in Table 1 are unchanged. However, the time-reversed single-particle eigenstates do change. This leads to different and additional quantum geometric quantities within the superfluid weight [22], potentially with a different low-temperature scaling, which can modify Table 1 in the TRS-breaking case.
The classification of superconducting states is done via the irreducible representations of the symmetry group of the system [34]. In particular, we need to distinguish between spin-singlets with total spin $S=0$ and spin-triplets with total spin $S=1$. The pair spin wave function of a spin-singlet is antisymmetric with respect to an exchange of the spin indices, i.e., the gap function can be expressed as
@@
where $\Delta_{\alpha\beta}(\mathbf{k})=\Delta_{\alpha\beta}(\mathbf{q}=0;\mathbf{k})$ and the function $f_{\alpha\beta}$ is determined by the symmetry group of the system. If $\Gamma$ denotes an irreducible representation of the group with dimension $d_{\Gamma}$, it is provided by
@@
where $\psi_{i}^{\Gamma}$ are the simplest basis functions that are even in $\mathbf{k}$ and respect the symmetry of the system and $\Delta_{\alpha\beta}^{i}$ are coefficients representing the order parameters of the superconductor. Similarly, a spin-triplet state with total spin $S=1$ has odd parity, i.e., the gap function can be expressed as
@@
$\Delta_{\alpha\beta}^{i}=\frac{U_{0}}{V}\sum_{\mathbf{k},n}\bar{\psi}_{i}^{\Gamma}(\mathbf{k})v_{n\alpha,+}(\mathbf{k})v^{\ast}_{n\beta,-}(\mathbf{k})[1-2n_{\mathrm{F}}(E_{\mathbf{k}n})]$ (8)
at $\mathbf{q}=0$. Here, the $2N_{\mathrm{B}}$ vectors $(v_{n,+},v_{n,-})$ represent the eigenvectors of the BdG Hamiltonian in the Nambu spinor basis. Moreover, we separate the overall temperature dependence of the gap function from its internal structure within the $d_{\Gamma}$-dimensional space of the irreducible representation. Concretely, we assume that all nonvanishing components share a single temperature-dependent amplitude $\Delta_{T}$ and differ only by fixed temperature-independent coefficients $\eta_{i}$, i.e., $\Delta^{i}_{\alpha\beta}\equiv\Delta_{T}\eta_{i}\delta_{\alpha\beta}$ for each $i=1,\ldots,d_{\Gamma}$, and we restrict $(\eta_{1},\ldots,\eta_{d_{\Gamma}})$ to a set of high-symmetry representative directions where each component is either absent or present with unit modulus. Examples include nematic, chiral, or cyclic states [36]. For a given choice of $(\eta_{1},\ldots,\eta_{d_{\Gamma}})$, the amplitude $\Delta_{T}$ and its temperature dependence is determined by the self-consistent equation,
$\Delta_{T}{\eta}_{i}=\frac{U_{0}}{2V}\sum_{\mathbf{k},n}\frac{\bar{\psi}_{i}^{\Gamma}(\mathbf{k})f(\mathbf{k})}{E_{\mathbf{k}n}}[1-2n_{\mathrm{F}}(E_{\mathbf{k}n})]\,$, (9)
for a fixed $i$ with $\eta_{i}\neq 0$, while the remaining equations provide a consistency check. In the following, we absorb the phase factors $\eta_{i}\neq 0$ into the basis functions.
The quasiparticle dispersion in a flat band can be approximated by $E_{\mathbf{k}}\approx|f(\mathbf{k})|$, i.e., the nodal structure of a flat-band superconductor is dictated by the gap function. Note that, since $f$ is an analytic function, once it vanishes on an arbitrary small open subset in momentum space, it necessarily needs to vanish on the whole domain according to the identity theorem, i.e., the gap function needs to be zero on the whole domain. Hence, in two-dimensional superconductors, a surface node (or Bogoliubov Fermi surface) cannot be realized. Let us shift and rotate the coordinate system such that the analytic function $f$ has a node at the origin and is not identically zero along the $k_{1}$ direction. The Weierstrass preparation theorem states that there exists a Weierstrass polynomial,
@@
As is clear from the factorization, the zero set of each linear factor with $\mathrm{Im}(\lambda_{i})=0$ produces a line node $k_{1}-\mathrm{Re}\,\lambda_{i}(k_{2})=0$ with shallowness $m_{i}$, while it becomes a point node of shallowness $m_{i}$ once $\mathrm{Im}(\lambda_{i})\neq 0$. To distinguish both types, we also write the factorization as
@@
exhibits a cusp at the origin. Nevertheless, the Newton-Puiseux theorem guarantees here that the functions $u_{i},v_{i},w_{i}$ admit convergent Puiseux expansions [54].
@@
Because the calculations become very complicated if one considers the general expression for the dispersion given in Eq. (15), in this work we restrict ourselves to the representative case with $v_{1,\ldots,J}\equiv 0$, $w_{1,\ldots,J}(k_{2})=k_{2}$, and
@@
where $m=qL+\sum_{j=1}^{J}p_{j}$ is the total radial order of vanishing, $L$ is associated to the number of straight nodal lines through the origin, and $q$ is the shallowness per line, i.e., $E_{\mathbf{k}}\sim\Delta_{T}|t|^{q}$ near any of the nodal lines if local coordinates with $t$ transverse to the line are taken. Moreover, to simplify the calculations, we assume $J=0$ if $L>0$, so that the pointlike enhancement at the origin is absent and $m=qL$ for the line nodes. This is enough for our physical application later.
@@ III.2 Line nodes
$D(E)=\frac{1}{4\pi^{2}}\int_{0}^{2\pi}\mathrm{d}\theta\int_{0}^{\Lambda}\mathrm{d}k\,k\,\delta(E-\Delta_{T}k^{m}|\cos(L\theta)|^{q})\,.$ (24)
@@
The incomplete beta function reads [16, §8.39]
@@
which, for $x=1$, corresponds to the usual beta function. Therefore, we find the DOS to be given by
@@
For $q/m=1/2$, we make use of the fact that $B_{x}(1/2,0)=2\mathrm{artanh}(\sqrt{x})$ [16, §9.12]. Since we have $\mathrm{artanh}(\sqrt{1-x^{2}})=\ln(2/x)+\mathcal{O}(x^{2})$ for $x\ll 1$, we find at leading order
@@
We proceed similarly for $q/m>1/2$. By using the power series of the hypergeometric function [16, §9.14], we find
@@
with $x=[E/(\Delta_{T}\Lambda^{m})]^{2/q}$ in our case. Thus, we obtain the DOS as
@@ IV.1 Order parameter
$d_{\Gamma}^{\prime}\Delta_{T}\approx\frac{U_{0}}{2V}\sum_{\mathbf{k}}\frac{|f(\mathbf{k})|^{2}}{\Delta_{T}E_{\mathbf{k}}}[1-2n_{\mathrm{F}}(E_{\mathbf{k}})]\,$, (34)
where we inserted $f(\mathbf{k})=\Delta_{T}\sum_{i=1}^{d_{\Gamma}}\eta_{i}\psi_{i}^{\Gamma}(\mathbf{k})$, cf. Eq. (4), and $d_{\Gamma}^{\prime}$ represents the number of nonzero coefficients $\eta_{i}\neq 0$. We further replace the momentum integral by an energy integral via Eq. (19). The gap equation becomes
$\Delta_{T}=\left[\frac{U_{0}}{2d_{\Gamma}^{\prime}}\int_{0}^{\infty}\mathrm{d}E\,D(E)E[1-2n_{\mathrm{F}}(E)]\right]^{1/2}\,.$ (35)
@@
$\Delta_{T}=\left[\frac{U_{0}\tilde{D}_{0}}{2d_{\Gamma}^{\prime}}\int_{0}^{\infty}\mathrm{d}E\,E^{\alpha}\ln^{\beta}(1/E)[1-2n_{\mathrm{F}}(E)]\right]^{1/(2+\alpha)}\,.$ (37)
@@
$\Delta_{0}-\Delta_{T}=\frac{U_{0}\tilde{D}_{0}}{d_{\Gamma}^{\prime}(\alpha+2)\Delta_{0}^{\alpha+1}}\int_{0}^{\infty}\mathrm{d}E\,\frac{E^{\alpha}\ln^{\beta}(1/E)}{e^{E/T}+1}\,.$ (38)
@@
$C_{\alpha}=\Gamma(\alpha+1)\eta(\alpha+1)\frac{U_{0}\tilde{D}_{0}}{d_{\Gamma}^{\prime}(\alpha+2)\Delta_{0}^{\alpha+1}}\,.$ (40)